OpenStudy (anonymous):

(2.25 x 10 squared(2) ) divided (2.5 x 10 to the 6th power)

OpenStudy (mathstudent55):

Is this the question? \( \dfrac{2.25 \times 10^2}{2.5 \times 10^6} \)

OpenStudy (anonymous):

yes @mathstudent55

OpenStudy (mathstudent55):

First, you need to divide 2.25 by 2.5 2.25/2.5 = 0.9 Then you divide the powers of 10: 10^2 / 10^6 = 10^-4 The quotient is 0.9 x 10^-4 This number is not in scientific notation. To put it in scientific notation you need to have a number from 1 to under 10 multiplied by a power of 10. 9 x 10^-5